# Estimating Treatment Effect Heterogeneity in Psychiatry: A Review and Tutorial with Causal Forests

### v1

@@ Abstract
Flexible machine learning tools are being used increasingly to estimate heterogeneous treatment effects. This paper gives an accessible tutorial demonstrating the use of the causal forest algorithm, available in the R package grf. We start with a brief non-technical overview of treatment effect estimation methods with a focus on estimation in observational studies, although similar methods can be used in experimental studies. We then discuss the logic of estimating heterogeneous effects using the extension of the random forest algorithm implemented in grf. Finally, we illustrate causal forest by conducting a secondary analysis on the extent to which individual differences in resilience to high combat stress can be measured among US Army soldiers deploying to Afghanistan based on information about these soldiers available prior to deployment. Throughout we illustrate simple and interpretable exercises for both model selection and evaluation, including targeting operator characteristics curves, Qini curves, area-under-the-curve summaries, and best linear projections. A replication script with simulated data is available at github.com/grf-labs/grf/tree/master/experiments/ijmpr.
Keywords: Causal inference, treatment heterogeneity, machine learning.
@@ 1 Introduction
Research on heterogeneity of treatment effects (HTE), as this variation in intervention effects has come to be known, has proliferated in recent years in psychiatry (e.g., Feczko et al.,, 2019; Allsopp et al.,, 2019; Kaiser et al.,, 2022; Cohen and DeRubeis,, 2018), and many other disciplines. The basic approach is to search for significant variation in treatment effects across subsamples. This is most easily done by estimating interactions between some hypothesized effect modifiers assessed prior to randomization and random assignment. When the interaction is non-zero, HTE is said to exist with respect to the specified variable—meaning that the treatment effect differs depending on the value of the variable.
@@
Numerous small clinical trials have reported HTE detections of this sort with respect to such fundamental baseline variables as patient age, sex, education, and symptom severity (e.g., Cuijpers et al.,, 2014; Driessen et al.,, 2010). But a question arises in trying to synthesize all this information when numerous significant interactions of this sort emerge, as meta-analyses show that they do (e.g., Maj et al.,, 2020). What is the best way to combine the information about all these specifiers into a single composite? A commonly used method is to apply the potential outcomes framework (Imbens and Rubin,, 2015) in analyzing the results of a moderated multiple regression model that includes multiple interactions. Counter-factual logic is used here to generate two predicted outcome scores from this model for each participant in the study, the first assuming that the participant was assigned to the intervention group and the second assuming that the participant was assigned to the control group (DeRubeis et al.,, 2014). The two within-person predicted outcome scores are then compared at the individual level and used as an estimate of the treatment effect for an individual based on the multivariate specifier profile of the individual. The distribution of these differences scores is then examined to investigate the pattern of significant individual differences in the treatment effect.
@@
A drawback of this approach is that in trying out many different regression models, applications of the procedure that do not account for multiple testing may lead to false positives (van Klaveren et al.,, 2019). This problem can be addressed, though, using data-driven machine learning methods to search for interactions with cross-validation to address the problem of over-fitting (VanderWeele et al.,, 2019; Wager and Athey,, 2018). The use of data-driven methods in this way also addresses the problem that the moderated regression approach only allows for a limited range of typically two-way interactions, whereas HTE can involve much more complex forms of interaction. Machine learning methods, in comparison, allow for these more complex forms of interaction to be discovered using nonparametric methods. A large body of work has made tremendous progress in adapting a variety of machine learning algorithms to estimate HTE. Examples include Athey and Imbens, (2016); Athey et al., (2019); Hahn et al., (2020); Kennedy, (2023); Künzel et al., (2019); Luedtke and van der Laan, (2016); Nie and Wager, (2021); Tian et al., (2014).
2 Estimating treatment effects in nonexperimental settings
@@
The assumption that our column of realized outcomes is consistent with the columns of potential outcomes, $Y_{i}=Y_{i}(W_{i})$, is called the Stable Unit Treatment Value Assumption (SUTVA). In some complicated social network settings, this assumption may not be viable. Consider for example if patient A is given a behavioral therapy and patient B is not. If patient A interacts with patient B through some social functioning, like a support group, then A’s newly acquired behavioral knowledge may influence B. Here, however, we rely on SUTVA and thus rule out scenarios like these by assuming no interference or spillover between patients.
@@
Even though for any single patient $i$ we never observe the difference $Y_{i}(1)-Y_{i}(0)$, we still have enough information to identify the average difference. If we collect a representative sample from our population and then randomly assign an intervention to some units in the sample while withholding it from others, then the simple difference in means (i.e., the prevalence among treated patients minus the mean prevalence among control patients) provides an unbiased estimate of the average treatment effect. The fact that the treatment assignment is randomized means that our hypothetical potential outcomes and treatment status are decoupled, and justifies this simple calculation.
The availability of such a randomized treatment assignment is typically constrained to randomized controlled trials. In many important settings, due to ethical and resource considerations, we don’t have access to a randomized intervention. A dataset where the intervention assignment is not randomized is often referred to as an observational dataset. As an example, via electronic health records, we could assemble a dataset that records an intervention, such as psychiatric hospitalization, together with an outcome, such as subsequent suicide, in a sample of patients presenting at an emergency department with suicidality. The decision whether to hospitalize such patients is clearly not randomized, but will rather depend on attending physician and hospital protocols, patient severity and perceived risk of suicidal behavior, available outpatient alternatives to hospitalization (e.g., day treatment programs), and availability of hospital beds. Some of these determinants of hospitalization are also likely to be determinants of subsequent suicidal behavior, which means that a simple comparison of the rates of suicidal behavior in, say, six months after emergency department presentation cannot be interpreted as providing evidence of the effects of hospitalization on these behaviors. At the same time, constructing a randomized controlled trial in a setting of this sort poses ethical challenges.
@@
which is the probability that a patient with confounder variables equal to $x$ is assigned the treatment. To estimate causal effects, a necessary condition is that the association between baseline predictors and intervention assignment is not so strong that no individuals have a nonzero probability of assignment to any of the interventions under study. This is the overlap assumption, which requires that the propensity score $e(x)$ is strictly above 0 and below 1 for every unit $x$. For example, if 100% of the suicidal patients presenting at an emergency department with severe thought disturbance were hospitalized, it would be impossible to use observational methods to evaluate the effect of hospitalization in preventing subsequent suicidal behaviors among such patients. It is only in the subset of patients whose probability of hospitalization is less than 100% and greater than 0% that data-driven estimates of treatment effects can be made. Assessing this requirement empirically translates into making sure that estimated propensity scores are not close to either zero or one, recognizing that estimation of treatment effects is impossible in the subset of patients whose scores are close to zero or one.
While the overlap assumption is testable, the direct conditional independence assumption under unconfoundedness is not, as it involves statements of independence between random variables. This is usually referred to as an identifying assumption and is fundamentally untestable. Instead, evaluation of identifying assumptions requires domain knowledge and subject-matter expertise to assess how credible they are in a particular application. In this tutorial, we are treating this first step as done or known, i.e., the underlying causal graph (Pearl,, 1995) is agreed upon. In the context of the example of psychiatric hospitalization as a predictor of subsequent suicidal behavior among suicidal emergency department patients, an example where this assumption is invoked is Ross et al., (2023).
To estimate average treatment effects in observational settings under unconfoundedness, we must account for how the different types of patients respond to and are assigned the treatment before making comparisons between average outcome scores. We can do this by estimating the propensity score. It is also possible to perform the confounding adjustment by estimating conditional outcome functions, or via a doubly robust combination of estimates of both the propensity and the conditional outcome function; see Ding, (2023), Hernan and Robins, (2023) and Wager, (2024) for recent textbook treatments of these methods.
@@ 2.1 From ATE to HTE
CATEs are averages of the individual-level treatment effects (ITEs), $\Delta_{i}=Y_{i}(1)-Y_{i}(0),$ which represent individual-level responses to treatment. If known, ITEs would directly represent a measure of HTE. Unfortunately, the ITE is not possible to estimate, given that we only get to observe one realized potential outcome for every patient $i$. An object that in granularity lies somewhere between the ITE and the ATE is the CATE,
@@
This more granular object has practical appeal since we can formulate it as a function, denoted by $\tau(\cdot)$, that takes as input some patient characteristics given by $x$, then maps it to a treatment effect, as given by the ATE for the subset of patients with this profile. An individualized treatment rule could be to treat those patients where the treatment effect, as measured by CATE, is sufficiently large. If only $x$ is known, the CATE thresholding is optimal (e.g., Bhattacharya and Dupas,, 2012).
@@ 3 The causal forest approach to estimating CATE
The challenge in targeting treatment heterogeneity that does not arise in the basic prediction setting is that scores on an outcome variable are observed for each participant, whereas individual-level treatment effects are not observed. As we saw previously, though, treatment effect estimation with nonexperimental data can be carried out by adjusting for baseline characteristics. We do this using a two-step process. In the first step, causal forests account for uneven treatment assignments via the propensity score $e(x)$, as well as baseline effects by estimating the conditional mean (marginalizing over treatment)
@@
In this first step, baseline effects are essentially stripped away via something known as an orthogonal construction (Chernozhukov et al.,, 2018; Robinson,, 1988). This construction works on the outcomes as well as on treatment indicators to generate residualized outcome $\widetilde{Y}_{i}=Y_{i}-m(X_{i})$ and treatment indicators $\widetilde{W}_{i}=W_{i}-e(X_{i})$ that have been adjusted for confounding effects. In the second step, $\widetilde{Y}_{i}$ and $\widetilde{W}_{i}$ are used to form treatment effects estimates, and a random forest is used to find predictor variable partitions where these treatment effects estimates exhibit heterogeneity.
@@
In the case where the analysis is applied to an experiment, where the probability of assignment to each intervention arm is known by construction, these probabilities can simply be supplied in the first step. However, in a nonexperimental situation in which we are attempting to approximate estimates of treatment effects by adjusting for nonrandom exposure respective to observed baseline covariates, the probabilities of treatment assignment are estimated in the first step with separate random forests.
@@
It is possible to modify this random forest framework to accommodate many other relevant statistical quantities and settings – thus the name “generalized” random forests–grf. For example, Athey et al., (2019) use this general framework to construct an instrumental forest that estimates HTE in settings with an instrumental variable, Cui et al., (2023) use this framework to construct a causal survival forest for estimating HTE with right-censored survival outcomes, and Friedberg et al., (2020) constructs a local linear forest. While these random forest algorithms are designed to directly target various causally relevant targets, we emphasize that one still needs to exercise caution when interpreting flexible non-parametric point estimates of these quantities. To give some intuition, machine learning models serve as great complements to traditional parametric statistical models as they allow us to be model-agnostic. This agnostic property, however, comes at the cost of introducing higher estimation uncertainty. So, our estimates of individual-level treatment effects come with considerable uncertainty, and in assessing the quality of these estimates, it is practical to move to a less granular summary measure. In the applied Section 4 we cover different approaches to assess if the individual-level estimates capture meaningful heterogeneity via calibration exercises. On a general note, the perspective we are taking is that machine learning tools can serve as effective algorithmic devices that target CATEs, which an analyst then can validate on held-out data, and transparently convey potentially interesting scientific findings regarding treatment heterogeneity.
@@
Finally, it is worth pointing out that the underlying identifying causal assumptions used in causal forests are the same as in classical parametric models, such as in the moderated regression approach described earlier. However, a main difference is that whereas classical parametric models (e.g., linear and logistic regressions) specify particular predictors and functional forms of their associations with the outcomes prior to estimation, causal machine learning approaches like causal forest use flexible data-driven forest-based constructions (grf also allow for more advanced use-cases where estimates of the first-stage inputs $e(x)$ and $m(x)$ can be supplied via separate machine learning models, such as for example gradient boosting or neural networks).

### v2

@@ Abstract
Flexible machine learning tools are increasingly used to estimate heterogeneous treatment effects. This paper gives an accessible tutorial demonstrating the use of the causal forest algorithm, available in the R package grf. We start with a brief non-technical overview of treatment effect estimation methods, focusing on estimation in observational studies; the same techniques can also be applied in experimental studies. We then discuss the logic of estimating heterogeneous effects using the extension of the random forest algorithm implemented in grf. Finally, we illustrate causal forest by conducting a secondary analysis on the extent to which individual differences in resilience to high combat stress can be measured among US Army soldiers deploying to Afghanistan based on information about these soldiers available prior to deployment. We illustrate simple and interpretable exercises for model selection and evaluation, including targeting operator characteristics curves, Qini curves, area-under-the-curve summaries, and best linear projections. A replication script with simulated data is available at github.com/grf-labs/grf/tree/master/experiments/ijmpr.
Keywords: causal inference, treatment heterogeneity, machine learning
@@ 1 Introduction
Research on heterogeneity of treatment effects (HTE), as this variation in intervention effects has come to be known, has proliferated in recent years in psychiatry (e.g., Feczko et al.,, 2019; Allsopp et al.,, 2019; Kaiser et al.,, 2022; Cohen and DeRubeis,, 2018), and many other disciplines. The basic approach is to search for significant variations in treatment effects across subsamples. This is most easily done by estimating interactions between some hypothesized effect modifiers assessed prior to randomization and random assignment. When the interaction is non-zero, HTE is said to exist with respect to the specified variable—meaning that the treatment effect differs depending on the value of the variable.
@@
Numerous small clinical trials have reported HTE detections of this sort with respect to such fundamental baseline variables as patient age, sex, education, and symptom severity (e.g., Cuijpers et al.,, 2014; Driessen et al.,, 2010). But a question arises in trying to synthesize all this information when numerous significant interactions of this sort emerge, as meta-analyses show that they do (e.g., Maj et al.,, 2020). What is the best way to combine the information about all these specifiers into a single composite? A commonly used method is to apply the potential outcomes framework (Imbens and Rubin,, 2015) in analyzing the results of a moderated multiple regression model that includes multiple interactions. Counterfactual logic is used here to generate two predicted outcome scores from this model for each participant in the study, the first assuming that the participant was assigned to the intervention group and the second assuming that the participant was assigned to the control group (DeRubeis et al.,, 2014). The two within-person predicted outcome scores are then compared at the individual level and used as an estimate of the treatment effect for an individual based on the multivariate specifier profile of the individual. The distribution of these differences scores is then examined to investigate the pattern of significant individual differences in the treatment effect.
@@
A drawback of this approach is that in trying out many different regression models, applications of the procedure that do not account for multiple testing may lead to false positives (van Klaveren et al.,, 2019). This problem can be addressed, though, using data-driven machine learning methods to search for interactions with cross-validation to address the problem of over-fitting (VanderWeele et al.,, 2019; Wager and Athey,, 2018). The use of data-driven methods in this way also addresses the problem that the moderated regression approach only allows for a limited range of typically two-way interactions, whereas HTE can involve much more complex forms of interaction. Machine learning methods, in comparison, allow for these more complex forms of interaction to be discovered using non-parametric methods. A large body of work has made tremendous progress in adapting a variety of machine learning algorithms to estimate HTE. Examples include Athey and Imbens, (2016); Athey et al., (2019); Hahn et al., (2020); Kennedy, (2023); Künzel et al., (2019); Luedtke and van der Laan, (2016); Nie and Wager, (2021); Tian et al., (2014).
2 Estimating treatment effects in non-experimental settings
@@
The assumption that our column of realized outcomes is consistent with the columns of potential outcomes, $Y_{i}=Y_{i}(W_{i})$, is called the Stable Unit Treatment Value Assumption (SUTVA). In some complicated social network settings, this assumption may not be viable. Consider, for example, if patient A is given behavioral therapy and patient B is not. If patient A interacts with patient B through some social functioning, like a support group, then A’s newly acquired behavioral knowledge may influence B. Here, however, we rely on SUTVA and thus rule out scenarios like these by assuming no interference or spillover between patients.
@@
Even though for any single patient $i$ we never observe the difference $Y_{i}(1)-Y_{i}(0)$, we still have enough information to identify the average difference. If we collect a representative sample from our population and then randomly assign an intervention to some units in the sample while withholding it from others, then the simple difference in means (i.e., the prevalence among treated patients minus the mean prevalence among control patients) provides an unbiased estimate of the average treatment effect. The fact that the treatment assignment is randomized means that our hypothetical potential outcomes and treatment status are decoupled, which justifies this simple calculation.
The availability of such a randomized treatment assignment is typically constrained to randomized controlled trials. In many important settings, due to ethical and resource considerations, we don’t have access to a randomized intervention. A dataset where the intervention assignment is not randomized is often referred to as an observational dataset. As an example, via electronic health records, we could assemble a dataset that records an intervention, such as psychiatric hospitalization, together with an outcome, such as subsequent suicide, in a sample of patients presenting at an emergency department with suicidality. The decision whether to hospitalize such patients is clearly not randomized but will rather depend on attending physician and hospital protocols, patient severity and perceived risk of suicidal behavior, available outpatient alternatives to hospitalization (e.g., day treatment programs), and availability of hospital beds. Some of these determinants of hospitalization are also likely to be determinants of subsequent suicidal behavior, which means that a simple comparison of the rates of suicidal behavior in, say, six months after emergency department presentation cannot be interpreted as providing evidence of the effects of hospitalization on these behaviors. At the same time, constructing a randomized controlled trial in a setting of this sort poses ethical challenges.
@@
which is the probability that a patient with confounder variables equal to $x$ is assigned the treatment. To estimate causal effects, a necessary condition is that the association between baseline predictors and intervention assignment is not so strong that no individuals have a non-zero probability of assignment to any of the interventions under study. This is the overlap assumption, which requires that the propensity score $e(x)$ is strictly above 0 and below 1 for every unit $x$. For example, if 100% of the suicidal patients presenting at an emergency department with severe thought disturbance were hospitalized, it would be impossible to use observational methods to evaluate the effect of hospitalization in preventing subsequent suicidal behaviors among such patients. It is only in the subset of patients whose probability of hospitalization is less than 100% and greater than 0% that data-driven estimates of treatment effects can be made. Assessing this requirement empirically translates into ensuring that estimated propensity scores are not close to either zero or one, recognizing that estimation of treatment effects is impossible in the subset of patients whose scores are close to zero or one.
While the overlap assumption is testable, the direct conditional independence assumption under unconfoundedness is not, as it involves statements of independence between random variables. This is usually referred to as an identifying assumption and is fundamentally untestable. Instead, the evaluation of identifying assumptions requires domain knowledge and subject-matter expertise to assess how credible they are in a particular application. In this tutorial, we are treating this first step as done or known, i.e., the underlying causal graph (Pearl,, 1995) is agreed upon. In the context of the example of psychiatric hospitalization as a predictor of subsequent suicidal behavior among suicidal emergency department patients, an example where this assumption is invoked is Ross et al., (2023).
To estimate average treatment effects in observational settings under unconfoundedness, we must account for how the different types of patients respond to and are assigned the treatment before making comparisons between average outcome scores. We can do this by estimating the propensity score. It is also possible to perform the confounding adjustment by estimating conditional outcome functions or via a doubly robust combination of estimates of both the propensity and the conditional outcome function. This is the approach grf implements in the function average_treatment_effect; see Ding, (2023), Hernan and Robins, (2023), and Wager, (2024) for recent textbook treatments of these methods.
@@ 2.1 From ATE to HTE
CATEs are averages of the individual-level treatment effects (ITEs), $\Delta_{i}=Y_{i}(1)-Y_{i}(0),$ which represent individual-level responses to treatment. If known, ITEs would directly represent a measure of HTE. Unfortunately, the ITE is not possible to estimate, given that we only get to observe one realized potential outcome for every patient $i$. An object that in granularity, lies somewhere between the ITE and the ATE is the CATE,
@@
This more granular object has practical appeal since we can formulate it as a function, denoted by $\tau(\cdot)$, that takes as input some patient characteristics given by $x$, then maps it to a treatment effect, as given by the ATE for the subset of patients with this profile. An individualized treatment rule could be to treat those patients where the treatment effect, as measured by CATE, is sufficiently large. If only $x$ is known, then CATE thresholding is optimal (e.g., Bhattacharya and Dupas,, 2012).
@@ 3 The causal forest approach to estimating CATE
(a) Centered outcomes ($\widetilde{Y}_{i}$) for control and treatment groups are plotted against $age$, a potential predictor variable.
(b) Estimated average treatment effects ($\hat{\tau}_{\text{left}}$ and $\hat{\tau}_{\text{right}}$) for subgroups split by $age$. The dashed vertical line indicates an optimal split point where the difference in treatment effect estimates between the left (blue) and right (red) samples is largest.
Figure 1: Stylized example of splitting to maximize treatment effect heterogeneity using centered outcomes and treatments.
The challenge in targeting treatment heterogeneity that does not arise in the basic prediction setting is that scores on an outcome variable are observed for each participant, whereas individual-level treatment effects are not observed. As we saw previously, though, treatment effect estimation with non-experimental data can be carried out by adjusting for baseline characteristics. We do this using a two-step process. In the first step, causal forests account for uneven treatment assignments via the propensity score $e(x)$, as well as baseline effects by estimating the conditional mean (marginalizing over treatment)
@@
In this first step, baseline effects are essentially stripped away via something known as an orthogonal construction (Chernozhukov et al.,, 2018; Robinson,, 1988). This construction works on the outcomes as well as on treatment indicators to generate residualized outcome $\widetilde{Y}_{i}=Y_{i}-m(X_{i})$ and treatment indicators $\widetilde{W}_{i}=W_{i}-e(X_{i})$ that have been adjusted for confounding effects. In the second step, $\widetilde{Y}_{i}$ and $\widetilde{W}_{i}$ are used to form treatment effects estimates, and a random forest is used to find predictor variable partitions where these treatment effects estimates exhibit heterogeneity. Figure 1 presents a stylized illustration of the underlying logic for this phase of the algorithm. For a candidate predictor variable, such as age, we seek a single axis-aligned split separating units into two groups with different average treatment effect estimates. Here, units with ages above 25 appear to benefit more, and a partition that immediately increases heterogeneity, as measured by the vertical distance between estimates in Figure 1(b), would be at this point. This exercise is repeated recursively for the left and right samples, considering all possible predictor variables (e.g., blood pressure) to partition the covariates $X_{i}$ into regions where treatment effects are expected to differ.
@@
In the case where the analysis is applied to an experiment, where the probability of assignment to each intervention arm is known by construction, these probabilities can simply be supplied in the first step. However, in a non-experimental situation in which we are attempting to approximate estimates of treatment effects by adjusting for nonrandom exposure respective to observed baseline covariates, the probabilities of treatment assignment are estimated in the first step with separate random forests.
@@
It is possible to modify this random forest framework to accommodate many other relevant statistical quantities and settings – thus the name “generalized” random forests–grf. For example, Athey et al., (2019) use this general framework to construct an instrumental forest that estimates HTE in settings with an instrumental variable, Cui et al., (2023) use this framework to construct a causal survival forest for estimating HTE with right-censored survival outcomes, and Friedberg et al., (2020) constructs a local linear forest. While these random forest algorithms are designed to directly target various causally relevant targets, we emphasize that one still needs to exercise caution when interpreting flexible non-parametric point estimates of these quantities. To give some intuition, machine learning models serve as great complements to traditional parametric statistical models as they allow us to be model-agnostic. This agnostic property, however, comes at the cost of introducing higher estimation uncertainty. So, our estimates of individual-level treatment effects come with considerable uncertainty, and in assessing the quality of these estimates, it is practical to move to a less granular summary measure. In the applied Section 4, we cover different approaches to assess if the individual-level estimates capture meaningful heterogeneity via calibration exercises. On a general note, the perspective we are taking is that machine learning tools can serve as effective algorithmic devices that target CATEs, which an analyst can then validate on held-out data and transparently convey potentially interesting scientific findings regarding treatment heterogeneity.
@@
Finally, it is worth pointing out that the underlying identifying causal assumptions used in causal forests are the same as in classical parametric models, such as in the moderated regression approach described earlier. However, a main difference is that whereas classical parametric models (e.g., linear and logistic regressions) specify particular predictors and functional forms of their associations with the outcomes prior to estimation, causal machine learning approaches like causal forest use flexible data-driven forest-based constructions (grf also allow for more advanced use-cases by allowing for user-specified estimates of the first-stage inputs $e(x)$ and $m(x)$, which can be derived from separate machine learning models, such as gradient boosting or neural networks, or through ensemble methods that combine multiple models (van der Laan et al.,, 2007)).
Remark 1.
In this context, an experimental setting is a special case of an observational setting with a known propensity score. The implication is that the only algorithmic difference between analyzing experimental or observational data is that in the experimental setting, we supply the known propensity score instead of estimating it (in causal_forest, this can be done via supplying the argument W.hat). The propensity score is then used the same way, including in downstream analyses, such as in calculating doubly robust average treatment effects discussed in Section 2 (statistical guarantees on average treatment effect estimation are then even stronger, see Wager, (2024, Corollary 3.3) for details).